OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = -1 /2 and b = 1

OpenStudy (anonymous):

y=mx+b where m=-1/2 abd b=1 now put everything together

OpenStudy (anonymous):

.so it will be y=-1/2x+1?

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

its the same problem as the one above

OpenStudy (anonymous):

oops sorry mistake i meant this one Input in standard form the equation of the given line. The line with m = -3/8 and passing through (5, -4)

OpenStudy (anonymous):

ok you have m=-3/8 now use this y-y1=m(x-x1) where x1=5 and y1=-4

OpenStudy (anonymous):

.i tried i couldnt get the answer

OpenStudy (anonymous):

farthest I got is to the y-4=-3/8x-15/8

OpenStudy (anonymous):